OpenStudy (anonymous):

Solve 6x=the square root of (24 +12x). Check for extraneous solutions.

OpenStudy (anonymous):

6x=sqrt(24+12x) (6x)^2 = 24+12x 36x^2 -12x -24 = 0 x= -2/3 or x= 1 if you substitute x=-2/3 on the main equation: 6x=sqrt(24+12x) 6*(-2/3)=sqrt(24+12(-2/3)) -4 = 4 if you substitute x=1 on the main equation: 6x=sqrt(24+12x) 6=sqrt(24+12) 6=6 so the only solution is x=1